### v2

# Tidal deformability in neutron stars from a microscopic point of view

Abstract
We present results for the tidal deformability in neutron stars, the tidal Love number $k_{2}$, and the effective deformability of a binary system. The microscopic equation of state for cold $\beta$-stable neutron matter is based upon high-precision two-neutron forces and includes the chiral three-neutron forces required at the chosen order. We review and motivate our choices for the high-density continuation of the microscopic equation of state. We discuss our predictions and observe that they are well within multimessenger constraints. In contrast, stiff equations of state that yield radii larger than about 13 km are ruled out by GW170817 constraints.
Keywords: Tidal deformability; neutron matter; neutron stars; chiral effective field theory
I Introduction
A fully microscopic equation of state (EoS) up to central densities of the most massive stars – potentially involving non-nucleonic degrees of freedom and phase transitions – is not within reach. Nevertheless, neutron stars are powerful natural laboratories for constraining theories of the nuclear EoS Abbott and others (2017); Abbott, B. P. and others (2017, 2018); Abbott and others (2019); Miller and others (2019, 2021) and studying matter under extreme conditions.
A mix of all fundamental forces existing in nature is present in a neutron star, which continues to motivate the large interdisciplinary interest in its properties and the intense effort to constrain those properties both observationally and experimentally, with various levels of statistical and systematic uncertainties, see e.g. Ref. Soumi et al. (2018). When working with microscopic predictions (as opposed to phenomenological models or statistical analyses), one must be mindful of the theory’s limitations and identify the best ways to extract and interpret information from the observational or terrestrial data.
With the first detection of gravitational waves (GW) on September 14, 2015 from LIGO, major accomplishments have been made within the fields of physics and astronomy Miller and Yunes (2019). Gravitational waves tied up one of the last loose ends in the theory of general relativity with remarkably high precision Cervantes-Cota and others (2016). Our present understanding of black holes and neutron stars also rely heavily on data collected from GW.
Just as black holes release immense gravitational waves, neutron stars are remarkable sources of these waves. Exotic processes are triggered within a neutron star’s core due to the enormous pressure, which can forcefully eject mass at relativistic speeds. Consequently, gravitational waves become vital means of measuring these fascinating astrophysical objects and the dynamic processes within them Lasky (2015). For gravitational waves to produce discernible signals, a crucial requirement is the presence of asymmetrical deformations in the neutron star. Alternative mechanisms, such as pulsar “glitches,” events where neutron stars undergo a sudden “jump” in spin frequency, or magnetar flares, where a neutron star with an exceptionally high magnetic dipole experiences magnetic decay, can also induce substantial gravitational waves Lasky (2015).
During a neutron star merger event, the gravitational field of one neutron star on the other induces tidal deformations, which are encoded in the late-inspiral gravitational wave signal. More precisely, when a neutron star is placed in the perturbing tidal gravitational field generated by its companion compact star, its shape becomes distorted, resulting in an induced quadrupole moment. The induced quadrupole moment of the neutron star will impact the energy of the system and increase the rate of emission of gravitational waves, with the result that the binary system emits more energy and evolves faster towards the collision and merger Flanagan and Hinderer (2008); Hinderer (2008). Clearly, compact binaries that are detectable through gravitational waves are natural systems to observe this type of interaction, which opens the way to finding additional constraints to the neutron star EoS. A stiffer EoS, where the pressure increases rapidly with density, yields stars with larger radii and larger tidal effects. Therefore, tidal distortion reflects the stiffness (softness) of the EoSs, with the inspiral proceeding more rapidly for stars with larger tidal deformability.
In this paper, we focus primarily on the tidal properties of neutron stars, which have a direct impact on the emitted gravitational wave signal, offering the opportunity of detecting the latter to constrain the former Faber and Rasio (2012); Baiotti and Rezzolla (2017); Baiotti (2019). This prospect was first realized in August of 2017, with the detection of gravitational waves generated by a coalescing neutron star binary. This groundbreaking event, known as GW170817 Abbott and others (2017), was made possible by the second-generation ground-based detectors LIGO Adams and others (2015) and Virgo Acernese and others (2015). We can expect future observations to detect the inspiral and coalescence of more compact binary systems.
As we review below, for a given neutron star mass, the tidal deformability is strongly correlated with the star radius and, thus, the star compactness, C, defined as M/R. Keeping in mind that accurate and precise neutron star radius measurements are challenging Khadkikar and et al. (2025), the importance of tight constraints on global neutron star properties that depend on M/R is self-evident. We also recall that the radius of the canonical neutron star correlates with the (model dependent) slope of the symmetry energy around saturation density.
Although this paper could be seen as an extension of Ref. Sammarruca and Ajagbonna (2025), we note that it is an important extension. The tidal deformability encodes the impact of the EoS on the phase evolution of the waveform, which we have not previously confronted with our EoS. Tidal deformation is yet another astrophysically observable macroscopic property of neutron stars that can be used to study their interiors. Currently, Bayesian statistical analyses are very popular in the nuclear physics community, as they should be. Our approach is not a statistical analysis, in that we make predictions remaining as close as possible to the chiral theory of nuclear forces, and follow new constraints as they become available. We believe that our findings in Ref. Sammarruca and Ajagbonna (2025) are valuable and should be tested against these independent data. Furthermore, while this paper focuses on a particular, fundamentally important physical property of neutron stars, we wish to place it within our ongoing microscopic studies of the nuclear equation of state, where we keep a consistent path from free-space nucleon-nucleon ($NN$) data to neutron-rich nuclei and to neutron stars, as illustrated in Fig. 1. “MBT” stands for many-body theory, while SNM, NM, and NS signify symmetric nuclear matter, neutron matter, and neutron-star matter, respectively.
Figure 1: (Color online) Schematic representation of our ab initio pipeline.
This paper is organized as follows. In Sec. II.1, we review briefly our theoretical ingredients, omitting details that have been published elsewhere. In Sec. III, we construct high-density continuations of the microscopic EoS, paying particular attention to the speed of sound in stellar matter. In Sec. IV.1, we outline the calculation of the tidal deformability, which can be easily incorporated in the solution scheme used for the well-known Tolman-Oppenheimer-Volkoff (TOV) differential equations for $\frac{dP}{dr}$ and $\frac{dm}{dr}$. Our results, along with other predictions and constraints, are discussed in Sec. IV.2. Our main takeaways are summarized in Sec. V.
II Theoretical tools
II.1 Basic concepts in chiral effective field theory
The first step in a development of an EFT is the identification of a “soft scale” and a “hard scale.” For this purpose, guidance can be found in the hadron spectrum, where a large separation exists between the mass of the pion and the mass of the vector meson $\rho$. It is therefore natural to identify the pion mass as the soft scale while the mass of the $\rho$ sets the hard scale, approximately 1 GeV. Moreover, since quarks and gluons are ineffective degrees of freedom in the low-energy regime, pions and nucleons are the appropriate degrees of freedom of the EFT. The connection between QCD and the EFT is established through the symmetries of low-energy QCD. At this point, we can write the most general Lagrangian consistent with those symmetries (and their breaking). Following the prescription of the theory as reviewed in Ref. Machleidt and Entem (2011), the QCD Lagrangian is given by:
$\mathcal{L}=\bar{q}(i\gamma^{\mu}\mathcal{D}_{\mu}-\mathcal{M})q-\frac{1}{4}\;\mathcal{G}_{\mu\nu,a}\mathcal{G}^{\mu\nu}_{a}\;,$ (1)
where ‘$q$’ is the quark field, $\mathcal{D}_{\mu}$ represents the gauge covariant derivative, $\mathcal{M}$ is the quark mass matrix, and $\mathcal{G}^{\mu\nu}_{a}$ is the gluon strength field tensor.
Chiral symmetry is conservation of “handedness,” and is an exact symmetry for massless particles. Chiral symmetry occurs in the limit of vanishing quark masses, which amounts to dropping the quark mass matrix term in the above Lagrangian. In fact, such term is responsible for explicit breaking of chiral symmetry, as can be seen from the following. The quark mass matrix,
$\mathcal{M}=\begin{pmatrix}m_{u}&0\\ 0&m_{d}\end{pmatrix}$ (2)
can be recast in terms of the identity matrix and the third Pauli spin matrix:
$\mathcal{M}=\frac{(m_{u}+m_{d})}{2}\begin{pmatrix}1&0\\ 0&1\end{pmatrix}+\frac{m_{u}-m_{d}}{2}\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}\;$. (3)
Clearly, the first term respects isospin symmetry while the second term vanishes if the masses of the “up” and “down” quarks are equal. Thus, the small difference in the quark masses breaks isospin symmetry. On the other hand, the expression above breaks chiral symmetry explicitly as a result of the non-zero quark masses. However, since the masses of the “uu” and “dd” quarks are very small compared to typical hadronic masses, explicit breaking of chiral symmetry is a small effect.
Next, we briefly address the spontaneous breaking of chiral symmetry, for which there is clear evidence in the hadron spectrum. The spontaneous breaking of a global (as opposed to local) symmetry is accompanied by the appearance of a so-called massless “Goldstone Boson.”The particle which fulfills these requirements is the pion, an isospin triplet pseudoscalar boson. The pion is light relatively to the other mesons in the hadron spectrum but not massless, which is due to the explicit chiral symmetry breaking from the non-vanishing quark masses.
Having identified pions and nucleons as the appropriate degrees of freedom, one can then proceed to construct the Lagrangian of the effective theory:
$\mathcal{L}_{eff}=\mathcal{L}_{\pi\pi}+\mathcal{L}_{\pi N}+\mathcal{L}_{NN}+$... (4)
This effective Lagrangian is expanded in the form of the “soft scale” over the “hard scale”, $\frac{Q}{\Lambda_{\chi}}$. $Q$ is of the order of the pion mass, whereas $\Lambda_{\chi}$ is the energy scale of chiral symmetry breaking, approximately 1 GeV.
Through a scheme known as power counting, the most important contributions to the effective Lagrangian are accounted for first, with increasing order resulting in consistently smaller terms. While the expansion itself is, of course, infinite, at each order we are assured that the number of terms to be retained is finite and well defined.
II.2 Chiral truncation error
Crucial to chiral EFT is the trucation error, defined below If observable $X$ is known at order $n$ and at order $n+1$, a reasonable estimate of the truncation error at order $n$ can be expressed as the difference between the value at order $n$ and the one at the next order:
$\Delta X_{n}=|X_{n+1}-X_{n}|\;,$ (5)
since this is a measure for what has been neglected at order $n$. To estimate the uncertainty at the highest order that we consider, we follow the prescription of Ref. Epelbaum et al. (2015). For an observable $X$ that depends on the typical momentum of the system under consideration, $p$, one defines $Q$ as the largest between $\frac{p}{\Lambda_{b}}$ and $\frac{m_{\pi}}{\Lambda_{b}}$, where $\Lambda_{b}$ is the breakdown scale of the chiral EFT, for which we assume 600 MeV. The uncertainty of the value of $X$ at N3LO is then given by:
$\Delta X=\max\{Q^{5}|X_{LO}|,Q^{3}|X_{LO}-X_{NLO}|,Q^{2}|X_{NLO}-X_{N^{2}LO}|,$
$Q|X_{N^{2}LO}-X_{N^{3}LO}|\}\;,$ (6)
where $p$ could be identified with the Fermi momentum at the density under consideration.
We close this section with a warning. While the predictions at N2LO are fully ab initio, a warning is in place for current N3LO calculations. As pointed out in Ref. Epelbaum et al. (2020), there is a problem with the regularized three-nucleon-force (3NF) at N3LO (and higher orders): in all present nuclear structure calculations. The N3LO 3NFs currently in use are regularized by a multiplicative regulator applied to the 3NF expressions as derived from dimensional regularization. This approach leads to violation of chiral symmetry at N3LO and destroys the consistency between two- and three-nucleon forces Epelbaum et al. (2020, 2023). Consequently, none of the current calculations that include 3NFs at N3LO (and beyond) can be considered truly ab initio. An appropriate symmetry-preserving regulator Epelbaum et al. (2020) should be applied to the 3NF at N3LO from Refs. Bernard et al. (2008, 2011). At the present time, reliable predictions exist only at N2LO, NLO, and LO. However, for the few fully ab initio calculations, the precision at N2LO is unsatisfactory. A first step towards deriving consistently regularized nuclear interactions in chiral EFT has been proposed in Refs. Krebs and Epelbaum (2023a, b). It requires the cutoff to be introduced already at the level of the effective Lagrangian. A path integral approach Krebs and Epelbaum (2023a) can then be applied to the regularized chiral Lagrangian to derive nuclear forces through the standard power counting of chiral EFT. We will come back to this point in our conclusions and outlook.
III The speed of sound in neutron star matter
III.1 Subluminar and non-conformal
The content of this section is based on Ref. Sammarruca and Ajagbonna (2025), where we discussed how causality and maximum-mass constraints pose considerable restrictions on the general features of the high-density EoS, if such continuation follows an ab intio portion.
Parametrizations of the high-density EoS in terms of the speed of sound can be used as an alternative to, or in combination with the popular polytropic extensions. Defining $\rho_{0}$, $\epsilon_{0}$, and $P_{0}$ as the baryon density, the energy density, and the pressure at threshold (the density at which the EoS parametrization is attached to the previous piece), we write Kanakis-Pegios et al. (2021); Tews et al. (2018), for $i\geq 1$,
$\rho_{i}=\rho_{i-1}+\Delta\rho\;,$ (7)
$\epsilon_{i}=\epsilon_{i-1}+\Delta\epsilon\;,$ (8)
and
$\Delta\epsilon=\Delta\rho\frac{\epsilon_{i-1}+P_{i-1}}{\rho_{i-1}}\;.$ (9)
A reasonable parametrization for the speed of sound is:
$\Big(\frac{v_{s}}{c}\Big)^{2}_{i}=1-c_{1}exp\Big[-\frac{(\rho_{i}-c_{2})^{2}}{w^{2}}\Big]\;,$ (10)
where the constants $c_{1}$ and $c_{2}$ are determined from continuity of the speed of sound and its derivative at the threshold density. We find this to be a practical choice for a function that should approach a limit asymptotically. The main point is to preserve causality, which could be done, for instance, by attaching additional polytropes with appropriate steepness so as to control maximum mass and causality.
Using the relation between the dimensionless speed of sound squared and $dP/d\epsilon$, we write the pressure above the threshold as
$P_{i}=\Big(\frac{v_{s}}{c}\Big)^{2}_{i-1}\Delta\epsilon+P_{i-1}\;,$ (11)
where we have This EoS continuation is obviously causal at all densities and allows for a maximum mass of 2.07 $M_{\odot}$.
We found Sammarruca and Ajagbonna (2025) that a better ansatz, for the purpose of achieving high maximum masses while respecting causality at any density, is to first continue the EoS with a relatively steep polytrope,
$P(\rho)=\alpha\rho^{\gamma}\;,$ (12)
and then follow with a parametrization as in Eqs. (1011), which will remain causal by construction. The matching densities are 0.277 $fm^{-3}$ and 0.563 $fm^{-3}$ Sammarruca and Ajagbonna (2025). We underlined that this scenario is fundamentally related to the softness of the chiral predictions. Therefore, the nature of the predictions at normal density have a far-reaching impact, which extends to densities up to a few times normal density.
That being said, piecewise polytropic extensions are quite general and remain a popular extension method. We found that, leveraging the flexibility offered by Eq.(12) over a limited high-density interval, would allow us to control the maximum mass, whereas the causality-preserving pressure-density relation is a more practical tool (than piecewise polytropes) to reach super-high densities.
Table 1: Some neutron star properties corresponding to the red and the blue M(R) relations shown on the left-hand side of Fig. 2. The Table is taken from Ref. Sammarruca and Ajagbonna (2025).
$\gamma$ chiral order $M_{max}/M_{\odot}$ $R_{M_{max}}(km)$ $\rho_{c}(fm^{-3})$ $R_{1.4}$ (km)
3.3 N2LO 2.19 10.39 1.09 11.84
N3LO 2.27 10.68 1.03 12.11
3.8 N2LO 2.43 11.06 0.93 12.09
N3LO 2.50 11.32 0.88 12.30
III.2 Subluminar and conformal
In the QCD limit of deconfined quarks in the presence of asymptotic freedom, quarks should behave like free fermions. Some perturbative QCD calculations Brandes et al. (2023) support the conformal limit, $(v_{s}/c)^{2}=1/3$. We observed that an EoS that’s subconformal at all densities (that is, Eq. (10) with the asymptotic limit replaced by 1/3), cannot generate sufficiently large masses. On the other hand, the speed of sound can be asymptotically conformal and non-monotone. A scenario where the speed of sound peaks around few to several times nuclear density and then falls back to approach the QCD limit for deconfined quark matter would signify some sort of phase transition with the conformal limit reached well beyond central densities of the heaviest observed neutron stars. The superconformality condition, $\Big(\frac{v_{s}}{c}\Big)^{2}>1/3$, satisfied on the average in neutron stars may have fundamental implications on the trace of the energy-momentum tensor at the center of rotating neutron stars Mendes et al. (2024).
Here, we will explore parametrizations where the speed of sound is consistent with the conformal limit as the density approaches infinity Tews et al. (2018). As an extension of Eq. (10), a natural choice might be to add a second, skewed, gaussian function, which approaches the asymptotic value of 1/3:
$\Big(\frac{v_{s}}{c}\Big)^{2}_{i}=\frac{1}{3}-c_{1}exp\Big[-\frac{(\rho_{i}-c_{2})^{2}}{w^{2}}\Big]+c_{3}exp\Big[-\frac{(\rho_{i}-c_{4})^{2}}{w_{2}^{2}}\Big]\Big[1+erf\Big(c_{5}\frac{(\rho_{i}-c_{4})}{w_{2}}\Big)\Big]\;.$ (13)
We solve for $c_{1}$ and $c_{2}$ as done above, matching Eq. (13) and its derivative to the values at threshold, within a chosen parameter space for the remaining three coefficients, $c_{3},c_{4},c_{5}$. The widths $w$ and $w_{2}$ are obtained by normalizing both gaussians in the standard way. Although this procedure can provide a very large number of possibilities, we observed that the roots, $c_{1}$ and $c_{2}$, converge easily within limited ranges for the free parameters: $c_{3}\approx[0,1]$, $c_{4}\approx[1,4]fm^{-3}$, $c_{5}\approx[-50,50]$. In fact, the roots tend to cluster around a few pairs of $c_{1},c_{2}$ values, weakly sensitive to the large variations in $c_{5}$, and only moderately sensitive to variations of $c_{3}$ and $c_{4}$. As a consequence, those parametrizations produce very similar M(R) results. A representative case is shown by the green curves on the left-hand side of Fig. 2, where we have chosen a polytropic index of 3.5 to ensure that the new curves would stand out in between the other two sets from. The corresponding curves for the dimensionless speed of sound are seen to approach the conformal limit on the right-hand side of Fig. 2. We also note that implementing the conformal limit at all densities would not allow sufficiently large masses. Maximum masses as large as those recently observed require a rapid rise of the pressure soon after the first matching point.
It is interesting to observe that, on the left panel of Fig. 2, the green curves are what one would expect from the non-conformal calculation and an adiabatic index that’s in between the other two. In contrast, the green predictions on the right are dramatically different from the red or blue predictions. This suggests that the maximum mass is only weakly sensitive to the super-high density behavior of the EoS.
Figure 2: Left: the red and blue M(R) curves are as described in Ref. Sammarruca and Ajagbonna (2025). Green curves: M(R) relations at N3LO (dashed) and at N2LO (solid) where the polytropic part of the EoS has adiabatic index equal to 3.5, and the last continuation is based on the conformal speed of sound parametrization, Eq. (13). Right: Dimensionless speed of sound squared corresponding to each of the cases displayed on the left. See text for additional explanation.
IV Tidal deformability
IV.1 General aspects
The tidal deformability is a parameter that quantifies how easily the star is deformed by an external tidal field. Therefore, a largetidal deformability signals a larger and less compact star which deforms easily.
A spherically symmetric star placed in a quadrupolar tidal field $\epsilon_{ij}$, the response to the perturbation is a quadrupole moment, $Q_{ij}$, which, to linear order in $\epsilon_{ij}$, can be written as
$Q_{ij}=-\lambda_{t}\epsilon_{ij}\;.$ (14)
In the Newtonian limit, the perturbing tidal field, $\epsilon_{ij}$, is the second spacial derivative of the external field, resulting in units of inverse length squared.The quadrupole moment has units of length cubed. We expect $\lambda_{t}$ to be proportional to the fifth power of the neutron star radius, $R$. Adopting traditional convention, we express the tidal deformability as
$\lambda_{t}=\frac{2}{3}k_{2}R^{5}\;,$ (15)
where $k_{2}$ is the gravitational Love number. Of greater astrophysical relevance is the dimensionless tidal deformability, defined as
$\Lambda=\frac{\lambda_{t}}{M^{5}}=\frac{2}{3}k_{2}C^{-5}\;,$ (16)
where C is the previously defined neutron star’s compactness. It is $\Lambda$ that actually governs the effects of tides on inspiral waveforms Van Oeveren and Friedman (2017).
Given a proposed equation of state, the tidal deformability of a neutron star of a certain mass can be obtained by computing the metric in the asymptotic regime using the Einstein equations, extracting the corresponding $r$-order terms, and taking their ratio. The ensuing calculation is described in detail in Ref. Hinderer (2008). The result is an expression for the gravitational Love number, $k_{2}=k_{2}(C,y_{R})$, with $y_{R}$ the solution at the star’s surface of the first order differential equation
$y^{\prime}(r)+(1/r)y^{2}(r)+(1/r)y(r)e^{\lambda(r)}{1+4\pi r^{2}[p(r)-\epsilon(r)]}+rQ(r)=0\;,$ (17)
In the equation above, the metric function for a spherical star is
$e^{\lambda}=[1-2m/r]^{-1}\;,$ (18)
and
$Q(r)=4\pi e^{\lambda(r)}[5\epsilon(r)+9p(r)+(\epsilon(r)+p(r))/c_{s}^{2}(r)-6e^{2\lambda(r)}/r^{2}-4e^{\lambda(r)}/r^{4}[m(r)+4\pi p(r)r^{3}]^{2}\;,$ (19)
where $m(r)$, $\epsilon(r)$, and $p(r)$ are the enclosed mass, the energy density, and the pressure at a radial distance $r$.
The Love number, $k_{2}$, is obtained in terms of $C$ and the value of $y(r)$ at $r=R$, $y_{R}$, from
$k_{2}(C,y_{R})=\frac{8}{5}C^{5}(1-2C)^{2}[2-y_{R}+2C(y_{R}-1)]\{2C[6-3y_{R}+3C(5y_{R}-8)]+$
$4C^{3}[13-11y_{R}+C(3y_{R}-2)+2C^{2}(1+y_{R})]+$
$3(1-2C)^{2}[2-y_{R}+2C(y_{R}-1)]ln(1-2C)\}^{-1}\;,$ (20)
and then used in Eq.( 16) to determine the tidal deformability.
Figure 3: Left: The dimensionless tidal deformability as a function of the neutron star mass (left) and as a function of the compactness (right). The pink dots and their uncertainties are taken from the Bayesian analysis of Ref. Lim and Holt (2019). See text for more details.
Figure 4: Left: The tidal Love number as a function of the neutron star mass (left) and as a function of the compactness (right).
IV.2 Results and discussion
IV.2.1 Tidal deformability and tidal Love number
[…]
Closely related to $\Lambda$ is the tidal Love number $k_{2}$, related to $\Lambda$ as in Eq. (16). This is shown in Fig. 4 as a function of the neutron star mass and compactness. Often, comparison between relativistic and Newtonian calculations or tabulations of deformation properties for polytropic EoS with different polytropic indices are given in terms of $k_{2}$ rather than $\Lambda$ Hinderer (2008). Our predicted values of $k_{2}$ are shown in Table 2 for a representative case.
Table 2: Compactness and tidal Love number as a function of the neutron star mass for one of the EoS used in Fig. 4.
EoS $M/M_{\odot}$ C $k_{2}$
N2LO $\gamma$=3.3 0.90 0.112 0.105
1.0 0.125 0.103
1.1 0.137 0.0972
1.2 0.150 0.0926
1.3 0.163 0.0853
1.4 0.175 0.0782
1.5 0.188 0.0722
1.6 0.202 0.0660
1.7 0.215 0.0579
1.8 0.230 0.0521
1.9 0.245 0.0428
2.0 0.261 0.0361
2.2 0.313 0.0161
Recent multimessanger constraints can be found in Ref. Huang (2025), obtained from various methodologies based on variants of the GW170817 data sets. Their most robust constraint is based on a combination of GW-based findings and EoS-independent X-ray constraints, and is reported as
$\Lambda_{1.4}=265.18^{+237.88}_{-104.38}\;\;\;\;\;\;\;\;R_{1.4}=11.53^{+0.89}_{-0.88}\;\mbox{km}\;,$ (21)
for the dimensionless tidal deformability and radius of the canonical mass neutron star. We underline that this EoS-independent constraint is yet another confirmation that the large radius extracted from the PREX neutron skin measurement Reed et al. (2021) is unrealistic.
We account for both chiral truncation error (cf. Sec. II.1), and the spreading due to different values of $\gamma$. Combining the two independent uncertainties in quadrature, we estimate our ranges for $\Lambda_{1.4}$ and $R_{1.4}$ at N2LO to be
$\Lambda_{1.4}=355.25\pm 85.64\;\;\;\;\;\;\;\;R_{1.4}=(11.99\pm 0.51)\;\mbox{km}\;,$ (22)
quite consistent with Eq. (21). We underline that the overall uncertainties in Eq. (22) are evaluated at N2LO, which, at this time, is a fully consistent order, see comments in Sec. II.1.
Other inferred upper limits on neutron star radii and tidal deformabilities from GW170817 rule out stiff equations of state that result in large radii and large tidal deformabilities. Recent estimates of the measurability of tidal effects and the ability of these observatories to constrain the EoS with signals from black hole-neutron star and binary neutron star systems can be found in Refs. Hotokezaka et al. (2016); Damour et al. (2012); Del Pozzo et al. (2013); Agathos and et al. (2015); Wade and et al. (2014) and references therein.
Next, we want to explore the sensitivity of $\Lambda$ to the high-density parametrization of the speed of sound discussed in Sec. III.2. For that purpose, we employ the four EoS used for Fig. 2, keeping the same color and pattern conventions. As for the M(R) results on the left side of Fig. 2, we see that the green curves are what we would expect from the non-conformal calculation and an adiabatic index that’s between the other two, while the speed of sound predictions with or without the conformal limit are dramatically different, see right side of Fig. 2. This seems to indicate that the tidal deformability is mostly sensitive to the EoS at medium to moderately high densities, a regime that “feels” the influence of the microscopic theory Sammarruca and Ajagbonna (2025), as discussed in Sec. III.1.
Figure 5: Tidal deformability as a function of mass (left) and compactness (right). The green curves are based on the EoS extended to the highest densities with a conformal parametrization of the speed of sound.
To make the above statement more quantitative, next we perform tests adopting different extension methods. Prior to exploring other options, our typical high-density extension consisted of piecewise polytropes with appropriately chosen adiabatic index. Equations of state unable to support a maximum mass of at least 2.01 M were discarded, and M(R) relations were cut at the central density where causality was violated. With these guidelines, we selected a few favorable combinations, given in Table 3.
Table 3: Maximum mass and $R_{1.4}$ generated by EoS continuations in terms of two polytropes with index $\Gamma_{1}$ and $\Gamma_{2}$, respectively.
EoS label $\Gamma_{1}$ $\Gamma_{2}$ $M_{max}/M_{\odot}$ $R_{1.4}$ (km)
A 3.1 2.7 2.10 12.00
B 3.1 2.8 2.12 12.00
C 3.2 2.7 2.15 12.06
D 3.2 2.8 2.17 12.06
E 3.3 2.7 2.19 12.11
In Table 4, we will compare Love numbers and $\Lambda_{1.4}$ for diverse high-density continuations.
Table 4: Maximum mass and $R_{1.4}$ generated by EoS continuations in terms of two polytropes with index $\Gamma_{1}$ and $\Gamma_{2}$, respectively.
EoS description $k_{2}$ $\Lambda_{1.4}$
A 0.083812 373.908325
B 0.083812 373.908325
C 0.085187 388.95812
D 0.085187 388.95812
E 0.086453 402.891215
N3LO, $\gamma$=3.3 0.086512 403.520110
$(v_{s}/c)^{2}$ $\rightarrow$ 1
N3LO, $\gamma$=3.8 0.091063 457.909912
$(v_{s}/c)^{2}$ $\rightarrow$ 1
N3LO, $\gamma$=3.3 0.086512 403.523460
($v_{s}/c)^{2}$ $\rightarrow$1/3
N3LO450, $\gamma$ = 3.8 0.091064 457.911549
($v_{s}/c)^{2}$ $\rightarrow$ 1/3
For the first group in Table 4, above the first horizontal line, the mild, if any, sensitivity to the high-density continuation is perhaps not surprising, since EoS A through E are not fundamentally different. The second group highlights the sensitivity to the polytropic index employed at medium-high densities. Finally, the third and last group, together with the second, highlights the sensitivity to the super-high density parametrization (for equal $\gamma$). We conclude that, in all of the above cases, differences are essentially negligible compared to the current uncertainties for these quantities. In particular, the last two entries in Table 4 demonstrate that the super-high density continuation has no impact on these observables.
IV.2.2 The mass-weighted tidal deformability
Here, we focus on our results for the effective tidal deformability, $\tilde{\Lambda}$, for a binary with masses (deformabilities) $M_{1}({\Lambda_{1}})$ and $M_{2}({\Lambda_{2}})$. First, we recall that $\Lambda_{1}$ or $\Lambda_{2}$ cannot be disentangled in the observed gravitational wave, and thus are not directly observable. Instead, it is possible to constrain the mass-weighted average (or effective) tidal deformability,
$\tilde{\Lambda}=\frac{16}{13}\frac{(M_{1}+12M_{2})M_{1}^{4}\Lambda_{1}+(M_{2}+12M_{1})M_{2}^{4}\Lambda_{2}}{(M_{1}+M_{2})^{5}}\;,$ (23)
with $M_{1}$ and $M_{2}$ the masses of the primary and the secondary neutron stars in the binary. $\tilde{\Lambda}$ determines the total effect of the tidal deformability on the phase evolution of the GW signal. Another observable that can be extracted from the evolution of the frequency in the observed GW signal is the chirp mass, defined as
$M_{c}=\frac{(M_{1}M_{2})^{3/5}}{(M_{1}+M_{2})^{1/5}}=M_{1}\frac{q^{3/5}}{(1+q)^{1/5}}\;,$ (24)
with $q=M_{1}/M_{2}$. The chirp mass is one of the most tightly constrained properties extracted from GW170817. Therefore, $M_{c}$ can be set equal to the observed value for the purpose of comparing predictions of $\tilde{\Lambda}$ with experimental data. The GW170817 constraint was stated as $M_{c}/M_{\odot}=1.188^{+0.004}_{-0.002}$ at the 90% confidence level Abbott and others (2017), from which the primary and secondary masses (in units of $M_{\odot}$), were inferred to be within the range (1.36, 1.60) and (1.17, 1.36), respectively, placing $q$ between 0.73 and 1. We exploited this information to calculate the values in Table 5. Starting with $M_{1}$ in the full range of masses inferred from GW170817, (1.17, 1.60), and imposing the very stringent $M_{c}$ constraint, we obtain $\tilde{\Lambda}$ for each value of $q$, see Table 5. The effective tidal deformability changes by about 4% across the $q$-variations. Essentially, $\tilde{\Lambda}$ is constrained at fixed mass.
Table 5: Effective tidal deformability under the GW170817 conditions.
$q=M_{2}/M_{1}$ $M_{1}/M_{\odot}$ $M_{2}/M_{\odot}$ $\tilde{\Lambda}$
0.73 (1.599,1.600) (1.167,1.168) (390.32, 388.61)
0.75 (1.577,1.584) (1.183,1.188) (393.76, 384.96)
0.80 (1.526, 1.532) (1.221,1.226) (399.19, 389.61[)
0.85 (1.479, 1.486) (1.257, 1.263) (401.30, 380.37)
0.90 (1.437, 1.443) (1.293, 1.299) (392.83, 382.89)
0.95 (1.398, 1.404) (1.328, 1.334) (396.11, 383.23)
1.0 (1.363, 1.369) (1.363, 1.369) (393.71, 383.98)
As we already pointed out when discussing $\Lambda_{1.4}$, one of the most physically interesting aspect of the GW170817 data is the opportunity to extract constraints on the radius of the canonical mass neutron star and thus discriminate among various EoS. In Fig. 6, the shaded area represents constraints from GW170817 and AT2017gfo Coughlin and others (2018), which confine $\tilde{\Lambda}$ to the interval (197, 720). The curve is a typical power law parametrization of $\tilde{\Lambda}$ as a function of $R_{1.4}$, see Ref. Koliogiannis et al. (2021) and references therein. Clearly, radii smaller than 10.8 km and larger than about 13 km are ruled out.
Figure 6: Effective deformability vs. the radius of the canonical mass neutron star. The shaded area represents constraints from GW170817 and AT2017gfo Coughlin and others (2018). The continuous curve is a parametrization of $\tilde{\Lambda}$ vs. $R_{1.4}$ Koliogiannis et al. (2021). Our predictions are indicated by the squares (q=0.73) and the crosses (q=1.0).
IV.2.3 Additional discussion
Our predictions are generally consistent with multimessenger constraints, as well as terrestrial laboratory constraints. For the purpose of rendering the paper more selfcontained, this short section will address recent terrestrial constraints.
For the symmetry energy, its density dependence, and the pressure in NM, vis a vis constraints from Ref. Lynch and Tsang (2022), the reader is referred to Ref. Sammarruca (2023). Reference Sammarruca (2024) presents a broad analysis of microscopic predictions versus empirical constraints for the neutron skin of 48Ca and 208Pb. In that work, we emphasize that the small skin of 48Ca extracted from CREX is quite consistent with microscopic predictions and available experimental constraints, see Fig.1 of Ref. Sammarruca (2023) and Ref. Sammarruca (2022).
The statistical analysis of Ref. Koehn and et al. (2025) employed 100,000 candidate EoS to obtain Bayesian posteriors on the symmetry energy parameters at normal density. The authors reanalyzed the results of PREX and CREX and concluded that neutron star radius posteriors can be statistically compatible with PREX and CREX. They discuss the question of how each constraint on the EoS is constructed from the observed data, and the various assumptions that could introduce biases. They comment on potential sources of systematic errors while applying a particular constraint, and discuss how different statistical interpretations of the data may affect the inferred results. The authors conclude “… the need for enhanced theoretical models in order to reduce the systematic uncertainties currently present and possibly accommodate both the PREX-II and CREX results.” We take the opportunity to add that theory improvements should never be directed at accomodating specific constraints.
Recently, new constraints Ciampi and et al. (2025) were obtained from the isospin transport ratio in 58,64Ni + 58,64Ni collisions at 32 MeV/nucleon, where the impact parameter dependence was experimentally measured. We understand that the BUU transport model employs effective nuclear energy density functionals that respect the present constraints from ab initio calculations. Figure 7 leaves little doubt on the “outlier” nature of the PREX-II result.
[…]
V Conclusions and outlook
As noted in Ref. Sammarruca and Ajagbonna (2025), the maximum-mass constraint moving to higher values, together with the causality requirement at any central density, poses significant restrictions on the high-density continuation of the EoS. Based on those considerations, we selected a small number of EoS differing in the chiral order and/or the adiabatic index of their polytropic portion, followed by a parametrization of the pressure vs. density that guarantees causality. We also used a parametrization where the speed of sound approaches the conformal limit at infinity and observed no significant impact on the M(R) relation or neutron star deformabilities.
We focused on the tidal deformability of neutron stars and related parameters, such as the tidal Love number. We addressed the effective tidal deformability of the binary and the implication of a very tight constraint on the chirp mass of the system. Measurement of the effective tidal deformability (at nearly fixed mass) can be used to extract constraints on the neutron star radius. This is very valuable, as precise neutron star radii would discriminate between equations of state better than any other single measurement. The upper limit on the effective tidal deformability from GW170817 excludes radii larger than 13.2 km, essentially independent of the assumed masses for the component stars. This rules out the findings from PREX II and any steep EoS that generates large radii. We underline the importance of EoS-independent measurements, which can help interpret divergent outcomes from terrestrial experiments that rely on EoS modeling assumptions. Most recent constraints Tang et al. (2015); Cui and et al. (2025); Ciampi and et al. (2025), from both neutron star observations and heavy ion collisions at Fermi energy, confirm the soft nature of the nuclear equation of state.
One of our main takeaways is that, while phenomenological extensions of the EoS are unavoidable, they can and must be controlled. Extensions of the (soft) chiral EFT EoS are generally consistent with both terrestrial and observational constraints. With the new generation of GW detectors and continuing progress toward ab initio nuclear theory, there is real hope to build a robust bridge connecting microscopic physics and telescope physics.
Acknowledgments
This work was supported by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences, under Award Number DE-FG02-03ER41270.
In Memoriam
F.S. dedicates this work to Rup, with everlasting love.